# Provable In-Context Learning of Nonlinear Regression with Transformers

### v1

@@ 3.2 One-Layer Transformer
$F(P;W^{K},W^{Q},W^{V})=W^{V}P\times\mathrm{softmax}\Big{(}(W^{K}P)^{T}W^{Q}P\Big{)}.$ (5)
@@
$W^{V}=\begin{pmatrix}0_{d\times d}&0_{d}\\ 0_{d}^{\top}&\nu\end{pmatrix},\quad W^{KQ}=\begin{pmatrix}Q&0_{d}\\ 0_{d}^{\top}&0\end{pmatrix},$
where $Q\in\mathbb{R}^{d\times d}$ is the trainable weight matrix. Such structured matrices separate out the impact of inputs and responses and have been justified to achieve the global or nearly global optimum for linear self-attention models in [18] and softmax attention models in [19], both for linear regression functions. In our setup, the choice of the last entry $W^{V}$ aligns with common practices in the theoretical analysis of deep networks, where the final linear layer is held constant to focus on understanding the behavior of hidden representations. The self-attention mapping becomes
@@
where we further let $\bar{X}=\begin{pmatrix}x_{1}&x_{2}&\cdots&x_{N}&x_{\mathrm{query}}\end{pmatrix}\in\mathbb{R}^{d\times N+1}$ and $\textbf{y}=\begin{pmatrix}y_{1}&y_{2}&\cdots&y_{N}\end{pmatrix}$$\in\mathbb{R}^{d\times N}$. The prediction $\hat{y}_{\mathrm{query}}$ corresponding to the query token $x_{\mathrm{query}}$ is given by the last entry of $F(P;Q)_{N+1}$, i.e., $\hat{y}_{\mathrm{query}}=F(P;Q)_{N+1}$. To train the attention model on the ICL problem introduced in Section 3.1, we minimize the following squared loss between the predicted and true responses:
@@
$\mathrm{vec}(Q)^{(t+1)}=\mathrm{vec}(Q)^{(t)}-\eta\nabla_{\mathrm{vec}(Q)}\mathcal{L}\big{(}\mathrm{vec}(Q)^{(t)}\big{)}.$ (8)
where $\eta$ is the learning rate. Based on this model setup and training procedure, we proceed to present our main theoretical results concerning ICL under nonlinear regression tasks.
@@ 4 Main Results of ICL Convergence
In this section, we present our main results on the convergence of the transformer model trained via GD for ICL with nonlinear regression functions. Interestingly, we observe that the flat and sharp regimes of the Lipschitz constant $L$ in (4) give rise to distinct convergence dynamics, with a threshold transition separating the two regimes based on the order of magnitude of $L$.
Recall from Section 3.1 that each token $x_{i}$ corresponds to a noisy version of a feature $v_{k}\in\mathbb{V}$ with probability $p_{k}=\Theta\left(\frac{1}{K}\right)$ for any $k\in[K]$. Let $P_{1:N}$ denote the collection of input tokens in $P$, i.e., $\{x_{i}\}_{i=1}^{N}$. We define the following concentration set of token sequences where each feature appears with approximately the expected frequency:
:=$\textstyle\mathcal{E}^{*}:=\Big{\{}P_{1:N}:|\mathcal{V}_{k}|\in\big{[}(p_{k}-\delta)N,(p_{k}+\delta)N\big{]}\mathrm{for}\ k\in[K]\Big{\}},$ (9)
@@
Building on the expression in (10), we characterize how the attention scores influence the prediction loss defined in (7) in the following lemma.
@@
Given constants $L,\Delta>0$, the prediction loss in (7) can be expressed as:
@@
The proof of Lemma˜1 is given in Appendix˜A. This expression reveals that the prediction loss $\mathcal{L}(P;Q)$ depends on the function Lipschitz constant $L$, the feature gap $\Delta$, and the attention score $\mathrm{Attn}_{k}^{(t)}$ associated with the true feature of the query token. For a given feature gap $\Delta$, different function Lipschitz constants $L$ can lead to distinct convergence behaviors of ICL. In the following theorem, we first provide the $\epsilon^{2}$-convergence of $\mathcal{L}(P;Q)$ in the flat $L$-regime, where $L$ is below a certain threshold.
@@
Suppose the function class in (4) satisfies $L\leq\Theta\left(\frac{1}{\Delta\delta}\right)$. Then, for any $0<\epsilon<1$ and under $N\geq\Theta(K^{3})$ and $K\geq\Theta(\frac{1}{\epsilon})$, with at most $T_{f}^{*}=\Theta(\frac{K\log(K\epsilon^{-1})}{\eta\delta^{2}L^{2}\Delta^{2}})$ iterations, we have $\mathcal{L}(P;Q)\leq\mathcal{O}(\epsilon^{2})$.
The proof of Theorem˜1 is given in Section˜C.3. Theorem˜1 indicates that $T_{f}^{*}$ decreases (i.e., the convergence is faster) as either the Lipschitz constant $L$ or the feature gap $\Delta$ increases. Intuitively, a larger function Lipschitz constant $L$ implies that the outputs $y_{i}$ and $y_{i^{\prime}}$ corresponding to different features in the prompt $P$ become more distinguishable, which facilitates faster in-context learning. Similarly, a larger feature gap $\Delta$ improves the separability among features, enabling the query token to more accurately attend to the relevant prompt tokens.
However, when $L$ exceeds a certain threshold, the caused sharp curvature induces large gradients on the attention weights. As a result, a smaller stepsize is required to stabilize convergence, leading to a convergence rate that differs from that in the flat regime. We next establish the $\epsilon^{2}$-convergence of $\mathcal{L}(\theta)$ in the sharp $L$ regime, where $L$ is above the threshold.
@@
Suppose the function class in (4) satisfies $L=\Omega(\frac{1}{\Delta\delta})$. Then for any $0<\epsilon<1$, under $N=\Omega(K^{3})$ and $K\geq\Theta(\frac{1}{\epsilon})$, with at most $T_{s}^{*}=\Theta(\frac{K\log(K\epsilon^{-1}L\Delta)}{\eta\epsilon\delta^{2}L^{2}\Delta^{2}})$ iterations, we have $\mathcal{L}(P;Q)=\mathcal{O}(\epsilon^{2})$.
The proof of Theorem˜2 is given in Section˜D.2. Theorem˜1 and Theorem˜2 together reveal an interesting phase transition phenomenon: the convergence dynamics are governed by how the function Lipschitz constant $L$ compares to a threshold of order $\Theta\left(\frac{1}{\Delta\delta}\right)$. In the flat curvature regime, where $L$ is below this threshold, convergence allows larger step sizes due to smaller gradients, resulting in a convergence rate of $\tilde{\Theta}\left(\frac{K}{\eta\delta^{2}L^{2}\Delta^{2}}\right)$. In contrast, in the sharp curvature regime, where $L$ is above the threshold, the large $L$ incurs large gradients which thus require smaller step sizes to stabilize convergence, yielding a convergence rate of $\tilde{\Theta}\left(\frac{K}{\eta\epsilon\delta^{2}L^{2}\Delta^{2}}\right)$. Comparing between the two convergence rates, for high-accuracy learning (i.e., small $\epsilon$), the flat regime may enjoy faster convergence. However, if $L$ is significantly larger than $\epsilon^{-1}$, the large $L$ results in much better distinguishability between features so that the attention matrix can align with the correct target much faster, and hence makes training much faster in the sharp regime.
@@
The proof of Proposition˜1 is given in Appendix˜E. This result follows directly from the loss expression in (11), where $\mathcal{L}(P;Q)=\Theta((1-\mathrm{Attn}_{k}^{(t)})^{2})$ when $x_{\mathrm{query}}=v_{k}$. Intuitively, as $\mathrm{Attn}_{k}^{(t)}$ approaches 111, the attention matrix effectively focuses predominantly on the prompt tokens that share the same underlying feature $v_{k}$. As a result, the predicted output $\hat{y}_{\mathrm{query}}$, given by (10), closely approximates the true value $f(v_{k})$, leading to high-accuracy predictions. Importantly, this does not require the function $f\in\mathcal{F}$ to have been seen during training, thereby enabling in-context learning at test time, i.e., the model, once trained to convergence, can produce correct predictions without any further fine-tuning.
@@ 5 Analysis of Convergence Dynamics
As established in Theorem˜1 and Theorem˜2, different regimes of the function Lipschitz constant $L$ lead to distinct convergence behaviors in ICL of nonlinear regression. In this section, we analyze the training dynamics under both regimes, highlighting how the Lipschitz constant $L$ influences the convergence rate in each case.
@@ 5.1 Gradients of Attention Weights
Based on the prediction output $\hat{y}_{\mathrm{query}}$ in (7), the attention scores $\mathrm{attn}_{i}$ play a critical role in determining the final prediction. To precisely characterize these attention scores for any $i\in[N]$, it is sufficient to characterize the training dynamics of the attention weights $q_{k,k^{\prime}}^{(t)}:=v_{k^{\prime}}^{\top}Q^{(t)}v_{k}$ for $k,k^{\prime}\in[K]$, which are initialized as $q^{(0)}_{k,k^{\prime}}=0$ for any $k,k^{\prime}\in[K]$. To simplify notations, we denote the $q_{k,k}^{(t)}$ as $q_{k}^{(t)}$ for $k^{\prime}=k$. According to the definition of the attention score $\mathrm{attn}_{i}^{(t)}$ in (10), when $x_{\mathrm{query}}^{(t)}=v_{k}$, the quantity $q_{k}^{(t)}$ measures how strongly the query token attends to the target feature $v_{k}$, while $q_{k,k^{\prime}}^{(t)}$ reflects the attention given to a different feature $v_{k^{\prime}}$ with $k^{\prime}\neq k$. To achieve the desired attention behavior, effective training should increase $q_{k}^{(t)}$ while suppressing $q_{k,k^{\prime}}^{(t)}$.
@@
For any $t\in[T]$, suppose $x_{\mathrm{query}}=v_{k}$. Then for any $k\in[K]$ and $k^{\prime}\in[K]$ with $k^{\prime}\neq k$, we obtain
@@
The proof of Lemma˜2 is given in Appendix˜B. From (12), we observe that $g_{k}^{(t)}$ is always non-negative, implying that the update $q_{k}^{(t)}$ increases over time. This growth continues until the attention score $\mathrm{Attn}_{k}^{(t)}$ approaches its convergence state near 111. However, $g_{k,k^{\prime}}^{(t)}$ in (13) is not necessarily positive and also depends on $\mathrm{Attn}_{k^{\prime}}^{(t)}$ associated with feature vector $v_{k^{\prime}}$. However, as $\mathrm{Attn}_{k}^{(t)}$ approaches 111, the residual term in (13) diminishes, and $g_{k,k^{\prime}}^{(t)}$ also converges toward zero, facilitating the overall convergence of the system.
@@ 5.2 Convergence Dynamics under Flat $L$-Regime
For ease of exposition, we consider the case where $x_{\mathrm{query}}=v_{k}$ in the following. Under the initialization of $q_{k}^{(0)}=q_{k,k^{\prime}}^{(0)}=0$, and by the definition of the attention score, we have $\mathrm{attn}^{(0)}_{n}=\frac{1}{N}$ for all $n\in[N]$, meaning that the transformer initially attends equally to all input tokens when computing the prediction for $x_{\mathrm{query}}$. We then leverage the task distribution in (4) and the gradient expressions in Lemma˜2 to analyze the learning dynamics of $q_{k}^{(t)}$ and $q_{k,k^{\prime}}^{(t)}$.
In the initial phase of training, the prediction $\hat{y}_{\mathrm{query}}$ is far from the ground truth $f(v_{k})$ due to the zero initialization of bilinear weights. According to (12), this results in a large positive gradient $g_{k}^{(t)}$, leading to a rapid increase in $q_{k}^{(t)}$. In contrast, the gradient $g_{k,k^{\prime}}^{(t)}$ may fluctuate in sign depending on the alignment of $f(v_{k})$ and $f(v_{k^{\prime}})$ at each step, causing $q_{k,k^{\prime}}^{(t)}$ to oscillate but decrease much more slowly. We formally characterize this phase below.
@@
The proof of Proposition˜2 is given in Section˜C.1. According to Proposition˜2, Phase I ends once $q_{k}^{(t)}$ becomes sufficiently large to reduce the prediction gap between the ICL output and the ground truth function value. After this point, both $g_{k}^{(t)}$ in (12) and $|g_{k,k^{\prime}}^{(t)}|$ in (13) decrease to smaller orders. During this phase, both $g_{k}^{(t)}$ and $|g_{k,k^{\prime}}^{(t)}|$ increase with $L$, as a larger function Lipschitz constant induces greater residual values in (12) and (13). Likewise, a larger feature gap $\Delta$ amplifies the value difference between $f(v_{k})$ and $f(v_{k^{\prime}})$ (by (4)), which in turn accelerates attention learning in Phase I. As a result, the duration $T_{f}^{1}$ decreases with both $L$ and $\Delta$.
However, at $t=T_{f}^{1}$, the prediction loss in (7) may still remain non-negligible. As a result, the attention score $\mathrm{Attn}_{k}^{(t)}$ requires a period of steady improvement after $t=T_{f}^{1}+1$. We now formalize this behavior in the second training phase in the following proposition.
@@
For any $t\in\{T_{f}^{1}+1,\cdots,T_{f}^{*}\}$ and $0<\epsilon<1$, where $T_{f}^{*}=\Theta(\frac{K\log(K\epsilon^{-1})}{\eta\delta^{2}L^{2}\Delta^{2}})$, $q_{k}^{(t)}$ continues to grow at a steady rate of $\Theta\left(\frac{\eta\delta^{2}L^{2}\Delta^{2}}{K}\right)$. Meanwhile, $q_{k,k^{\prime}}^{(t)}$ oscillates at a slower rate of $\mathcal{O}(\frac{\eta\delta^{2}L^{2}\Delta^{2}}{K^{2}})$ and exhibits an overall decreasing trend. At $t=T_{f}^{*}+1$, if $L$ satisfies $L\leq\Theta(\frac{1}{\Delta\delta})$ in (4), we have $\mathrm{Attn}_{k}^{(T_{f}^{*}+1)}=\Omega(\frac{1}{1+\epsilon\delta})$.
The proof of Proposition˜3 is given in Section˜C.2. According to Proposition˜3, if the function Lipschitz constant is sufficiently small such that $L\leq\Theta(\frac{1}{\Delta\delta})$, then by the end of Phase II, the attention score satisfies $1-\mathrm{Attn}_{k}^{(T_{f}^{*}+1)}=\mathcal{O}\left(\frac{\epsilon\delta}{1+\epsilon\delta}\right)=\mathcal{O}(\epsilon)$, indicating that the transformer has reached convergence.

### v2

@@ 3.2 One-Layer Transformer
$F(P;W^{K},W^{Q},W^{V})=W^{V}P\times\mathrm{softmax}\Big((W^{K}P)^{\top}W^{Q}P\Big).$ (5)
@@
$W^{V}=\begin{pmatrix}0_{d\times d}&0_{d}\\ 0_{d}^{\top}&1\end{pmatrix},\quad W^{KQ}=\begin{pmatrix}Q&0_{d}\\ 0_{d}^{\top}&0\end{pmatrix}$,
where $Q\in\mathbb{R}^{d\times d}$ is the trainable weight matrix. These simplifications, while not capturing the full complexity of deep, multi-head models, are standard in the theoretical literature and serve two crucial purposes. First, they allow for a tractable analysis that isolates the core dynamics of the softmax attention mechanism, which is central to ICL. Second, this setup follows a line of foundational work that has successfully used similar models to derive key insights into ICL for linear regression. By building on this framework, we can directly investigate the impact of nonlinearity. We acknowledge that extending our analysis to multi-layer and multi-head settings is an important avenue for future work. Such structured matrices separate out the impact of inputs and responses and have been justified to achieve the global or nearly global optimum for both linear and softmax attention models in [18] and [19]. Then the self-attention mapping becomes
@@
where we further let $\bar{X}=\begin{pmatrix}x_{1}&x_{2}&\cdots&x_{N}&x_{\mathrm{query}}\end{pmatrix}\in\mathbb{R}^{d\times N+1}$ and $\textbf{y}=\begin{pmatrix}y_{1}&y_{2}&\cdots&y_{N}\end{pmatrix}$$\in\mathbb{R}^{d\times N}$. The prediction $\hat{y}_{\mathrm{query}}$ corresponding to $x_{\mathrm{query}}$ is given by the last entry of $F(P;Q)_{N+1}$, i.e., $\hat{y}_{\mathrm{query}}=F(P;Q)_{N+1}$. To train the attention model on the ICL problem introduced in Section 3.1, we minimize the following squared loss between the predicted and true responses:
@@
$\mathrm{vec}(Q)^{(t+1)}=\mathrm{vec}(Q)^{(t)}-\eta\nabla_{\mathrm{vec}(Q)}\mathcal{L}\big(P;\mathrm{vec}(Q)^{(t)}\big).$ (8)
where $\eta>0$ is the learning rate. Based on this model setup and training procedure, we proceed to present our main theoretical results concerning ICL under nonlinear regression tasks.
@@ 4 Main Results of ICL Convergence
Our analysis proceeds by decomposing the loss function into interpretable quantities that directly reflect cluster separation and the Lipschitz constant $L$ in Eq.˜4. Interestingly, we observe that the flat and sharp regimes of $L$ give rise to distinct convergence dynamics, with a threshold transition separating the two regimes based on the order of magnitude of $L$. In the flat regime (small $L$), the gradient on $Q$ matrices remains small, leading to slow but steady concentration of attention weights. In the sharp regime (large $L$), we show a rapid growth phase where the query–key inner products amplify differences between clusters before settling into a slow fine-tuning phase. For the two regimes, we provide explicit convergence-time bounds and characterize the phase transition. Compared with prior analyses restricted to linear tasks or orthogonal features, our framework extends to general nonlinear Lipschitz tasks and explains qualitatively different dynamics observed in practice.
Recall from Section 3.1 that each token $x_{i}$ corresponds to a noisy version of a feature $v_{k}\in\mathbb{V}$ with probability $p_{k}=\Theta\left(\frac{1}{K}\right)$ for any $k\in[K]$. Let $P_{1:N}$ denote the collection of input tokens in $P$, i.e., $\{x_{i}\}_{i=1}^{N}$, and denote $\mathcal{V}_{k}\subset[N]$ as the index set for input tokens, such that $x_{i}=v_{k}$ for $i\in\mathcal{V}_{k}$. We define the following concentration set of token sequences where each feature appears with approximately the expected frequency:
$\textstyle\mathcal{E}^{*}:=\Big\{P_{1:N}:|\mathcal{V}_{k}|\in\big[(p_{k}-\delta)N,(p_{k}+\delta)N\big]\mathrm{for}\ k\in[K]\Big\},$ (9)
@@
Building on the expression in Eq.˜10, we characterize how the attention scores influence the prediction loss defined in Eq.˜7 in the following lemma.
@@
Given constants $L,\Delta>0$, the prediction loss in Eq.˜7 can be expressed as:
@@
The proof of Lemma˜1 is given in Appendix B. This expression reveals that loss $\mathcal{L}(P;Q)$ depends on the Lipschitz constant $L$, the feature gap $\Delta$, and the attention score $\mathrm{Attn}_{k}^{(t)}$ associated with the true feature of the query token. The dependence on $L$, which captures the function’s curvature, distinguishes our analysis from prior theoretical work on ICL (e.g., [19, 27, 23]). For a given feature gap $\Delta$, different function Lipschitz constants $L$ can lead to distinct convergence behaviors of ICL. In the following theorem, we first provide the $\epsilon^{2}$-convergence of $\mathcal{L}(P;Q)$ in the flat $L$-regime, where $L$ is below a certain threshold.
@@
Suppose the function class in Eq.˜4 satisfies $L\leq\Theta\left(\frac{1}{\Delta\delta}\right)$. Then, for any $0<\epsilon<1$ and under $N\geq\Theta(K^{3})$ and $K\geq\Theta(\frac{1}{\epsilon})$, with at most $T_{f}^{*}=\Theta(\frac{K\log(K\epsilon^{-1})}{\eta\delta^{2}L^{2}\Delta^{2}})$ iterations, we have $\mathcal{L}(P;Q)\leq\mathcal{O}(\epsilon^{2})$.
The proof of Theorem˜1 is given in Appendix D.3. Theorem˜1 indicates that $T_{f}^{*}$ decreases (i.e., the convergence is faster) as either the Lipschitz constant $L$ or the feature gap $\Delta$ increases. Intuitively, a larger function Lipschitz constant $L$ implies that the outputs $y_{i}$ and $y_{i^{\prime}}$ corresponding to different features in the prompt $P$ become more distinguishable, which facilitates faster in-context learning. Similarly, a larger feature gap $\Delta$ improves the separability among features, enabling the query token to more accurately attend to the relevant prompt tokens.
However, when $L$ exceeds a certain threshold, the caused sharp curvature induces large gradients on the attention weights. As a result, a smaller stepsize is required to stabilize convergence, leading to a convergence rate that differs from that in the flat regime. We next establish the $\epsilon^{2}$-convergence of $\mathcal{L}(P;Q)$ in the sharp $L$ regime, where $L$ is above the threshold.
@@
Suppose the function class in Eq.˜4 satisfies $L=\Omega(\frac{1}{\Delta\delta})$. Then for any $0<\epsilon<1$, under $N=\Omega(K^{3})$ and $K\geq\Theta(\frac{1}{\epsilon})$, with at most $T_{s}^{*}=\Theta(\frac{K\log(K\epsilon^{-1}L\Delta)}{\eta\epsilon\delta^{2}L^{2}\Delta^{2}})$ iterations, we have $\mathcal{L}(P;Q)=\mathcal{O}(\epsilon^{2})$.
The proof of Theorem˜2 is given in Appendix E.2. Theorem˜1 and Theorem˜2 together reveal an interesting phase transition phenomenon: the convergence dynamics are governed by how the function Lipschitz constant $L$ compares to a threshold of order $\Theta\left(\frac{1}{\Delta\delta}\right)$. In the flat curvature regime, where $L$ is below this threshold, convergence allows larger step sizes due to smaller gradients, resulting in a convergence rate of $\tilde{\Theta}\left(\frac{K}{\eta\delta^{2}L^{2}\Delta^{2}}\right)$. In contrast, in the sharp curvature regime, where $L$ is above the threshold, the large $L$ incurs large gradients which thus require smaller step sizes to stabilize convergence, yielding a convergence rate of $\tilde{\Theta}\left(\frac{K}{\eta\epsilon\delta^{2}L^{2}\Delta^{2}}\right)$. Comparing between the two convergence rates, for high-accuracy learning (i.e., small $\epsilon$), the flat regime may enjoy faster convergence. However, if $L$ is significantly larger than $\epsilon^{-1}$, the large $L$ results in much better distinguishability between features so that the attention matrix can align with the correct target much faster, and hence makes training much faster in the sharp regime.
@@
The proof of Proposition˜1 is given in Appendix F. This result follows directly from the loss expression in Eq.˜11, where $\mathcal{L}(P;Q)=\Theta((1-\mathrm{Attn}_{k}^{(t)})^{2})$ when $x_{\mathrm{query}}=v_{k}$. Intuitively, as $\mathrm{Attn}_{k}^{(t)}$ approaches $1$, the attention matrix effectively focuses predominantly on the tokens that share the same feature $v_{k}$. As a result, the predicted output $\hat{y}_{\mathrm{query}}$, given by Eq.˜10, closely approximates the true value $f(v_{k})$, leading to high-accuracy predictions. Importantly, this does not require the function $f\in\mathcal{F}$ to have been seen during training, enabling in-context learning at test time [22], i.e., the model, once trained to convergence, can produce correct predictions without any further fine-tuning.
@@ 5 Analysis of Convergence Dynamics
As established in Theorem˜1 and Theorem˜2, different regimes of the Lipschitz constant $L$ lead to distinct convergence behaviors in ICL. In this section, we analyze the training dynamics under both regimes, highlighting how the Lipschitz constant $L$ influences the convergence rate in each case.
@@ 5.1 Gradients of Attention Weights
Based on the prediction output $\hat{y}_{\mathrm{query}}$ in Eq.˜7, the attention scores $\mathrm{attn}_{i}$ play a critical role in determining the final prediction. To precisely characterize these attention scores for any $i\in[N]$, it is sufficient to characterize the training dynamics of the attention weights $q_{k,k^{\prime}}^{(t)}:=v_{k^{\prime}}^{\top}Q^{(t)}v_{k}$ for $k,k^{\prime}\in[K]$, which are initialized as $q^{(0)}_{k,k^{\prime}}=0$ for any $k,k^{\prime}\in[K]$. To simplify notations, we denote the $q_{k,k}^{(t)}$ as $q_{k}^{(t)}$ for $k^{\prime}=k$. According to the definition of the attention score $\mathrm{attn}_{i}^{(t)}$ in Eq.˜10, when $x_{\mathrm{query}}^{(t)}=v_{k}$, the quantity $q_{k}^{(t)}$ measures how strongly the query token attends to the target feature $v_{k}$, while $q_{k,k^{\prime}}^{(t)}$ reflects the attention given to a different feature $v_{k^{\prime}}$ with $k^{\prime}\neq k$. To achieve the desired attention behavior, effective training should increase $q_{k}^{(t)}$ while suppressing $q_{k,k^{\prime}}^{(t)}$.
@@
For any $t\in[T]$, suppose $x_{\mathrm{query}}=v_{k}$. Then for any $k,k^{\prime}\in[K]$ with $k^{\prime}\neq k$, we obtain
@@
The proof of Lemma˜2 is given in Appendix C. From Eq.˜12, we observe that $g_{k}^{(t)}$ is always non-negative, implying that the update $q_{k}^{(t)}$ increases over time. This growth continues until the attention score $\mathrm{Attn}_{k}^{(t)}$ approaches its convergence state near $1$. However, $g_{k,k^{\prime}}^{(t)}$ in Eq.˜13 is not necessarily positive and also depends on $\mathrm{Attn}_{k^{\prime}}^{(t)}$ associated with feature vector $v_{k^{\prime}}$. However, as $\mathrm{Attn}_{k}^{(t)}$ approaches $1$, the residual term in Eq.˜13 diminishes, and $g_{k,k^{\prime}}^{(t)}$ also converges toward zero, facilitating the overall convergence of the system.
@@ 5.2 Convergence Dynamics under Flat $L$-Regime
For ease of exposition, we consider the case where $x_{\mathrm{query}}=v_{k}$ in the following. Under the initialization of $q_{k}^{(0)}=q_{k,k^{\prime}}^{(0)}=0$, and by the definition of the attention score, we have $\mathrm{attn}^{(0)}_{n}=\frac{1}{N}$ for all $n\in[N]$, meaning that the transformer initially attends equally to all input tokens when computing the prediction for $x_{\mathrm{query}}$. We then leverage the task distribution in Eq.˜4 and the gradient expressions in Lemma˜2 to analyze the learning dynamics of $q_{k}^{(t)}$ and $q_{k,k^{\prime}}^{(t)}$.
In the initial phase of training, the prediction $\hat{y}_{\mathrm{query}}$ is far from the ground truth $f(v_{k})$ due to the zero initialization of bilinear weights. According to Eq.˜12, this results in a large positive gradient $g_{k}^{(t)}$, leading to a rapid increase in $q_{k}^{(t)}$. In contrast, the gradient $g_{k,k^{\prime}}^{(t)}$ may fluctuate in sign depending on the alignment of $f(v_{k})$ and $f(v_{k^{\prime}})$ at each step, causing $q_{k,k^{\prime}}^{(t)}$ to oscillate but decrease much more slowly. We formally characterize this phase below.
@@
The proof of Proposition˜2 is given in Appendix D.1. According to Proposition˜2, Phase I ends once $q_{k}^{(t)}$ becomes sufficiently large to reduce the prediction gap between the ICL output and the ground truth function value. After this point, both $g_{k}^{(t)}$ in Eq.˜12 and $|g_{k,k^{\prime}}^{(t)}|$ in Eq.˜13 decrease to smaller orders. During this phase, both $g_{k}^{(t)}$ and $|g_{k,k^{\prime}}^{(t)}|$ increase with $L$, as a larger function Lipschitz constant induces greater residual values in Eq.˜12 and Eq.˜13. Likewise, a larger feature gap $\Delta$ amplifies the value difference between $f(v_{k})$ and $f(v_{k^{\prime}})$ (by Eq.˜4), which in turn accelerates attention learning in Phase I. As a result, the duration $T_{f}^{1}$ decreases with both $L$ and $\Delta$.
However, at $t=T_{f}^{1}$, the prediction loss in Eq.˜7 may still remain non-negligible. As a result, the attention score $\mathrm{Attn}_{k}^{(t)}$ requires a period of steady improvement after $t=T_{f}^{1}+1$. We now formalize this behavior in the second training phase in the following proposition.
@@
For any $t\in\{T_{f}^{1}+1,\cdots,T_{f}^{*}\}$ and $0<\epsilon<1$, where $T_{f}^{*}=\Theta(\frac{K\log(K\epsilon^{-1})}{\eta\delta^{2}L^{2}\Delta^{2}})$, $q_{k}^{(t)}$ continues to grow at a steady rate of $\Theta\left(\frac{\eta\delta^{2}L^{2}\Delta^{2}}{K}\right)$. Meanwhile, $q_{k,k^{\prime}}^{(t)}$ oscillates at a slower rate of $\mathcal{O}(\frac{\eta\delta^{2}L^{2}\Delta^{2}}{K^{2}})$ and exhibits an overall decreasing trend. At $t=T_{f}^{*}+1$, if $L$ satisfies $L\leq\Theta(\frac{1}{\Delta\delta})$ in Eq.˜4, we have $\mathrm{Attn}_{k}^{(T_{f}^{*}+1)}=\Omega(\frac{1}{1+\epsilon\delta})$.
The proof of Proposition˜3 is given in Appendix D.2. According to Proposition˜3, if the Lipschitz constant is sufficiently small such that $L\leq\Theta(\frac{1}{\Delta\delta})$, then by the end of Phase II, the attention score satisfies $1-\mathrm{Attn}_{k}^{(T_{f}^{*}+1)}=\mathcal{O}\left(\frac{\epsilon\delta}{1+\epsilon\delta}\right)=\mathcal{O}(\epsilon)$, indicating that the transformer has converged.